OpenStudy (rahulmr):

let f to be a differentiable function defined for all real x, where f(x) >=0 for all x[0,a]. If \[\int\limits_{0}^{a}f(x) dx=a\], then \[2\int\limits\limits_{0}^{5a}(f(\frac{ x }{ 5 })+3) dx\] must be equal to Can anyone help me in this question. Don't know what to do. Thanks.

OpenStudy (samigupta8):

Did you start by substitution?

OpenStudy (rahulmr):

because it said that f(x)=, so i did this\[2\int\limits\limits\limits_{0}^{5a}(f(\frac{ a }{ 5 })+3) dx\] and i got 10a*(a+3) but my teacher said that the answer must be linear.

OpenStudy (samigupta8):

Okay! Let me start then.. First of all, you should have commenced like substitution of x/5 as t. Then what will your next step? Can you write it?

ganeshie8 (ganeshie8):

Rahul x is not a May I know why you're replacing x by a ?

OpenStudy (samigupta8):

Good point!

ganeshie8 (ganeshie8):

And yeah you should start as suggested by sami. It gives you a nice integral to work with

ganeshie8 (ganeshie8):

\(\int\limits_{0}^{a}f(x) dx=a\) This expression means the area under the graph of f(x) over the interval (0, a) is 'a'

OpenStudy (rahulmr):

ok thanks. i'll try.

OpenStudy (rahulmr):

@samigupta8 i didn't get what you meant by x/5 as t. Can you explain it a bit more.

OpenStudy (samigupta8):

You know substituition?

OpenStudy (samigupta8):

replcae dx by dt and also chamge the limits accordingly.

OpenStudy (samigupta8):

Replace*

OpenStudy (rahulmr):

\[2\int\limits\limits_{0}^{5t}(f(t)+3) dt\] would it be this

OpenStudy (samigupta8):

Limits are 0-t.

OpenStudy (samigupta8):

Not 0-5t

OpenStudy (rahulmr):

ok. like that \[2\int\limits\limits_{0}^{t}(f(t)+3) dt\]

OpenStudy (samigupta8):

Yep.

OpenStudy (rahulmr):

now should i integrate ?

OpenStudy (samigupta8):

Go ahead!

OpenStudy (rahulmr):

what is f(t). do i need to know that

OpenStudy (samigupta8):

Okay! Let me start wity basics then.

OpenStudy (rahulmr):

i got \[2(\int\limits\limits_{0}^{t}f(t) dt+3t)\]

OpenStudy (samigupta8):

If you have integral of f(x)dx as a then what can you say for integral of f(t)dt ?

OpenStudy (rahulmr):

t

OpenStudy (samigupta8):

Ah! No..

OpenStudy (samigupta8):

It is a.

OpenStudy (rahulmr):

sorry. i don't know then

OpenStudy (samigupta8):

Okay! You can easily get it if you replace t by x .

OpenStudy (samigupta8):

Then you get dt as dx. So we could see the integration for f(t)dt reduces to ?

OpenStudy (samigupta8):

Did you get it?

OpenStudy (rahulmr):

does it reduces to a

OpenStudy (samigupta8):

Yes obviously!

OpenStudy (rahulmr):

so now ??

OpenStudy (samigupta8):

Integrate ..

OpenStudy (rahulmr):

i got (a+3)t. is this correct

OpenStudy (samigupta8):

I got 2(a+3t)

OpenStudy (rahulmr):

oh yeah. i forgot to multiply with 2

OpenStudy (samigupta8):

You even multiplied the term a with t. Which i didn't

OpenStudy (rahulmr):

and now what do we do ??

OpenStudy (samigupta8):

Now solve this .. Put the limits from 0-5a

OpenStudy (samigupta8):

Don't forget that limit is for t !

OpenStudy (rahulmr):

I don't have enough time to do this question. I have to do other questions as well. so can you show me how to do this question. Thanks.

OpenStudy (samigupta8):

I have guided you how to solve this problem. I don't know how should i go about giving you the ans. Maybe @ganeshie8 devise a better way to help you out. Sorry , i can't help more into this.

OpenStudy (rahulmr):

Ok. thanks for the help.

OpenStudy (rahulmr):

@ganeshie8. can you help me in this question

OpenStudy (rahulmr):

@welshfella can you help me in this question

OpenStudy (welshfella):

I'm very rusty at this stuff I'm afraid (long time since I did it) I think you plug in t = 0 and t=5a into 2 (a + 3t) so we get 2[ a + 3(5a) - ( a + 3(0)]

OpenStudy (welshfella):

which simplifies to a linear function

OpenStudy (welshfella):

sorry i cant be of more help

OpenStudy (rahulmr):

thanks for the help. i will write 30a as my answer.

OpenStudy (samigupta8):

It has to be 32a.